Well i have to solve the following system of equations by substitution
0.5x-2y=1
-3y+x=4
I understand that I must get of the the equations to equal to either x or y so i can substitute.
i'll try the second one
-3y+x=4
x=4-3y is that right? if so, what to do next?

Answer by stanbon(48559)

0.5x-2y=1
-3y+x=4
I understand that I must get of the the equations to equal to either x or y so i can substitute.
i'll try the second one
-3y+x=4
x=4-3y is that
---------------
Substitute that into 5x-2y=1 to solve for "y":
0.5(4-3y)- 2y = 1
2 - (3/2)y - 2y = 1
-7/2 y = -1
y = 2/7
--------------------
Substitute into x = 4-3y
x = 4-3(2/7)
x = 4-6/7
x = 22/7
-----------------------------
Checking:
0.5x-2y = 1
0.5(22/7) -2(2/7) = 1
11/7 - 4/7 = 1
7/7 = 1
1 = 1
===========
Cheers,
Stan H.